OpenStudy (anonymous):

Solve: x^2(x^4 + 16) = 8x^4

OpenStudy (anonymous):

take x^2 =t u get t(t^2 + 16) =8t^2

OpenStudy (anonymous):

how did you get t?

OpenStudy (anonymous):

just take x^2=y now y(y^2 +16)=8y^2 y{y^2-8y+16}=0 y{(y-4)^2}=0 y=0 or y-4=0 means y=0 or 4 means x^2=0 or 4 means x=0 or +2 or -2

OpenStudy (anonymous):

thanks s but i still dont get this problem

OpenStudy (anonymous):

wts wrong?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

x = 0; x = -2; x = 4

OpenStudy (anonymous):

i keep getting that